# VeriLLM: A Lightweight Framework for Publicly Verifiable Decentralized Inference

### v3

@@ 5.2 Registration
  1. 1.
    the scheduler’s VRF verification public key, enabling third parties to validate the pseudorandom outputs used in node assignment and hidden-state sampling [micali1999verifiable];
  2. 2.
@@
The VRF key ensures that all sampling and scheduling randomness remains publicly auditable and resistant to precomputation or selective bias [buterin2014next, wood2014ethereum].
@@
This mechanism parallels the delayed-withdrawal model of accountable distributed systems [castro1999practical, luu2016smartpool], guaranteeing that both schedulers and workers remain economically bound to their past behavior.
5.3 Scheduling Scheme
@@
Verifiable randomness and auditability. Each scheduling decision is deterministically derived from a publicly verifiable random beacon instantiated via a verifiable random function (VRF) [micali1999verifiable]. The Scheduler’s VRF verification key is registered on-chain, and every scheduling round includes the tuple:
$(\texttt{job\_id},\texttt{VRF\_out},\texttt{VRF\_proof}),$ (6)
where VRF_out encodes the pseudorandom seed used for node assignment and hidden-state sampling, and VRF_proof allows any observer or contract to independently validate correctness. This construction guarantees that role assignment and verification checkpoints are unpredictable yet verifiable, preventing any party from biasing job allocation.
@@
$G_{m,\zeta}=\{n\in\mathcal{N}|\texttt{model}(n)=m,~\texttt{slice}(n)=\zeta\},$ (7)
@@
$\mathcal{P}(m)=\langle G_{m,1},G_{m,2},\dots,G_{m,L}\rangle.$ (8)
@@
$T_{\text{total}}=\sum_{\zeta=1}^{L}\big(T_{\zeta}+T_{\text{compute}}^{(\zeta)}\big),$ (9)
@@
5.4 Task Scheduling and Hidden-State Management
@@
$h\leftarrow H(\texttt{req}),\qquad r_{i}\leftarrow\mathrm{VRF}(sk_{\mathsf{sch}},h,i,L_{i}),$ (10)
@@
$\mathrm{VRF.Verify}(vk_{\mathsf{sch}},\langle h,i,L_{i}\rangle,r_{i},\pi_{i})=1.$ (11)
@@
$n^{\mathrm{inf}}_{i}=G_{m,i}[\sigma_{i}(1)],\qquad\{n^{\mathrm{ver}}_{i,1},\dots,n^{\mathrm{ver}}_{i,k}\}=G_{m,i}[\sigma_{i}(2..k{+}1)].$ (12)
@@
$S^{(t)}_{2}=g_{1}\big(\mathrm{Embed}(\mathbf{y}_{<t})\big).$ (13)
Each subsequent segment $i=2,\dots,L$ receives the preceding output $S^{(t)}_{i}$, executes its corresponding model layers, and produces:
$S^{(t)}_{i+1}=g_{i}(S^{(t)}_{i})$. (14)
At the final stage, the tail segment computes the logits and predicts the next token:
$y_{t+1}=\arg\max f_{\mathrm{LM}}\big(S^{(t)}_{L+1}\big)$. (15)
@@
Hidden-State Logging and Commitment. To enable verifiable post-hoc auditing, each Inferencer logs its intermediate hidden states for every token and commits to them using cryptographic proofs. For each token index $t$, an Inferencer serializes its output tensor $S^{(t)}_{i}\in\mathbb{R}^{d_{k}\times d_{v}}$, constructs a Merkle tree $\mathsf{Merkle}(S^{(t)}_{i})$ over all scalar entries, and submits the Merkle root $\gamma^{(t)}_{i}$ accompanied by a bound signature:
$\texttt{Sig}_{\mathrm{inf}}\big(h,i,t,\gamma^{(t)}_{i}\big)$. (16)
To prevent tampering or substitution during relay, the Scheduler also signs the forwarded root:
$\texttt{Sig}_{\mathrm{sch}}\big(h,i,t,\gamma^{(t)}_{i}\big)$, (17)
ensuring accountability for both the producer and the forwarder. This dual-signature commitment model provides end-to-end traceability: any discrepancy between a segment’s output and the forwarded record can be attributed with cryptographic certainty. Because verifiers only require Merkle openings for sampled positions, the bandwidth and on-chain verification costs scale sublinearly with tensor size. The integration of deterministic serialization, Merkle-based commitment [merkle1989certified], and chained signatures provides tamper-evident provenance for all hidden states without introducing measurable latency into the inference path.
Overall, this architecture ensures that every stage of distributed inference remains cryptographically auditable while preserving throughput and latency comparable to optimized large-model serving systems [rajbhandari2022deep, shoeybi2019megatron].
5.5 Inference Scheme
Figure 2: Overview of the decentralized inference architecture. Each segment hosts identical Transformer layers, and hidden states are relayed across the pipeline under cryptographic commitments.
The inference process executes online, autoregressive generation over a decentralized, pipelined fabric of homogeneous node groups. Each group hosts an identical copy of a contiguous Transformer block, including weights, normalization statistics, and kernel configurations [shoeybi2019megatron, rajbhandari2022deep]. For each incoming task, the Scheduler selects eligible groups and, through a verifiable random function (VRF) [micali1999verifiable], assigns one Inferencer per group to execute the live forward pass. Verifier roles are simultaneously drawn from the same VRF distribution but remain inactive during this phase. Throughout generation, the Scheduler relays and logs hidden states across segment boundaries, appending cryptographic commitments and signatures to construct an auditable trace that serves as the foundation for subsequent verification [zhang2020delphi, chen2024zkml].
@@ 5.5 Inference Scheme
$S^{(t)}_{1}$ $\leftarrow\texttt{Embed}(x_{1:t}),$ (18)
$S^{(t)}_{i+1}$ $\leftarrow F_{i}\big(S^{(t)}_{i};W_{i}\big),\quad i=1,\dots,L,$ (19)
$y_{t}$ $\leftarrow\texttt{Decode}\big(S^{(t)}_{L+1}\big)$, (20)
where $S^{(t)}_{i}$ is the hidden state entering segment $i$, $F_{i}$ denotes the segment’s forward operator, $W_{i}$ its parameters, and Decode the final output head combined with the sampling rule. The resulting pipeline achieves token-serial yet segment-parallel execution, synchronized by the Scheduler’s relay of hidden states [rajbhandari2022deep].
Scheduler-Driven Role Assignment. For each inference task, the Scheduler enumerates all node groups capable of collectively serving the model and assigns roles in a publicly verifiable manner. Using a VRF bound to the task identifier [micali1999verifiable], it deterministically selects one Inferencer per group, while keeping the randomness unpredictable before the VRF proof is revealed. This mechanism ensures that role assignment cannot be biased or precomputed and that every selection remains verifiable through the VRF proof posted on-chain [bonneau2015sok, qu2025zkgpt].
@@
  2. 2.
    computes a cryptographic commitment to $S^{(t)}_{i+1}$ (e.g., a Merkle root over a deterministic serialization [merkle1989certified]) and signs the root with its node key; and
  3. 3.
@@
The Scheduler verifies each signature against the on-chain registry (optionally through a zero-knowledge membership proof [chen2024zkml, sun2024zkllm] to preserve role anonymity), appends its own relay signature, persists the state to stable storage, and forwards the package to the next segment’s Inferencer. To minimize latency, each receiving node can begin its computation immediately while performing signature verification asynchronously; if any proof later fails, the Scheduler aborts the affected pipeline and invalidates dependent results.
@@
All node public keys are registered under an on-chain Merkle accumulator, enabling compact membership proofs [micali1999verifiable, bonneau2015sok]. Instead of forwarding raw signatures, the Scheduler may alternatively issue a zero-knowledge proof attesting that it holds a valid signature from some registered key on the advertised root—thereby maintaining verifiability without revealing node identity [sun2024zkllm, qu2025zkgpt]. Each forwarded artifact is re-signed by the Scheduler to guarantee origin authenticity and replay protection. This layered commitment and signature discipline ensures that every hop in the relay chain is both tamper-evident and cryptographically attributable.
Autoregressive Token Emission and Feedback. The terminal segment ($i=L$) outputs logits or the final hidden state for token $t$, from which the next token $y_{t}$ is sampled according to the decoding policy [brown2020language]. The Scheduler collects $y_{t}$ and feeds it back into the pipeline to form $x_{1:t+1}$, advancing the process to step t+1t{+}1. The loop continues until the model emits an end-of-sequence token $\langle\texttt{EOS}\rangle$ or another termination condition is met.
Task-Type Indistinguishability and Robustness. The communication protocol, message format, and API surface used by Inferencers during inference are identical to those used by Verifiers in the subsequent verification phase [yuan2022decentralized]. Consequently, no node can distinguish in real time whether it is performing live inference or participating in a verification probe. This indistinguishability eliminates the opportunity for strategic deviation and simplifies implementation symmetry across both phases. Combined with VRF-driven role assignment, signed commitments, and auditable relays, the inference scheme produces a verifiable, tamper-evident execution trace with negligible online overhead.
5.6 Verification Scheme
Figure 3: Overview of the decentralized verification architecture. Each verifier independently reconstructs its segment’s outputs via a full-sequence prefill, comparing committed and recomputed states.
During autoregressive inference, an LLM executes two phases: a prefill pass—where the prompt tokens are processed in parallel to populate the key–value (KV) cache—and a decode phase—where tokens are generated sequentially with causal dependencies on all previous outputs [brown2020language, shoeybi2019megatron]. The decoding phase dominates latency and is inherently serial, leading to limited GPU utilization and memory-bound performance [rajbhandari2022deep]. However, verification differs fundamentally: once the full output sequence $y_{1:T}$ is known, the verifier is no longer constrained by sequential dependencies. Instead, it can recompute all token representations simultaneously under the same causal mask, yielding a single parallel pass equivalent to prefill over the concatenated prompt–output sequence [sun2024zkllm, zhang2020delphi].
@@ 5.6 Verification Scheme
$x=[\texttt{prompt}\|y_{1:T}],$ (21)
and let $\mathcal{S}=[S^{\text{prompt}}\|S^{1}\|\dots\|S^{T}]$ denote the concatenated hidden states across all positions. Given the causal structure of the Transformer, the verifier can execute a single forward pass to reconstruct all hidden states $\{S^{(t)}_{i}\}_{t=1}^{T}$ for every segment $i\in[1,L]$ without performing incremental token decoding [vaswani2017attention]. This transformation replaces $T$ serial decode iterations with a single batched forward computation, eliminating redundant KV-cache updates and reducing cross-node communication. In practice, this yields significantly improved GPU utilization, especially on architectures with high Tensor Core throughput [rajbhandari2022deep].
@@
$T_{\text{verify}}\approx\frac{1}{L}T_{\text{prefill}}(\text{full model}),$ (22)
yielding near-linear scaling in the number of participating verifier groups [rajbhandari2022deep, zhang2020delphi].
Figure 4: Verification workflow. Each verifier independently executes a full-sequence prefill over its model segment and commits its output root to the blockchain for later consistency checks.
Prefill Computation. After the inference phase concludes, the Scheduler initiates segment-local verification by distributing the required input states to each verifier [qu2025zkgpt]. For segment $i>1$, the Scheduler sends the sequence $\{S_{i}^{(t)}\}_{t=1}^{T}$—the hidden states produced by the upstream inferencer for all tokens. For the first segment ($i=1$), the Scheduler instead provides the full tokenized input $\texttt{prompt}\|y_{1:T}]$, corresponding to the raw embedding input of the model. Upon receipt, each verifier executes a full-sequence prefill limited to its hosted layers:
$S_{i+1}^{(t)}=F_{i}(S_{i}^{(t)};W_{i}),\quad\forall t\in[1,T],$ (23)
@@
Importantly, the communication interface, data serialization format, and RPC protocol used during verification are identical to those used in live inference [yuan2022decentralized]. This design choice enforces task-type indistinguishability: from the perspective of any node, it is impossible to discern whether the current request corresponds to live inference or a verification probe [bonneau2015sok]. Such indistinguishability eliminates mode-specific behavior and prevents nodes from selectively deviating or shortcutting computation when acting as verifiers. Upon completion, each verifier must return its computed outputs $\{S^{(t)}_{i+1}\}_{t=1}^{T}$ through the same interface as during inference, producing an auditable trace indistinguishable in structure and timing.
Commitment Submission. Once the local recomputation completes, the Scheduler transitions verifiers into the commit phase. Each verifier constructs a tamper-evident cryptographic commitment to its final outputs—specifically, the hidden state corresponding to the last token ($t=T$) at its segment’s output boundary. This commitment is instantiated as a Merkle root [merkle1989certified]:
$\texttt{root}^{(T)}_{i}=\texttt{MerkleRoot}\big(S^{(T)}_{i+1}\big)$, (24)
@@
$\sigma^{(T)}_{i}=\texttt{Sig}(\texttt{sk}_{i},\texttt{root}^{(T)}_{i})$, (25)
@@
and publishes the commitment on-chain through the designated verification contract [chen2024zkml, sun2024zkllm].
@@
This on-chain record provides a permanent anchor for later verification steps. In subsequent sampling or dispute phases, verifiers may be required to disclose specific slices or elements of the hidden state $\{S^{(T)}_{i+1}\}$, accompanied by Merkle inclusion proofs that can be efficiently checked against the posted root [bonneau2015sok, micali1999verifiable]. This approach ensures both compactness and auditability: the entire verification process involves only lightweight commitments and selective disclosures rather than full tensor uploads, keeping blockchain overhead minimal while preserving cryptographic integrity.
@@
In summary, the verification scheme leverages the known output sequence and the pre-recorded intermediate states to convert sequential autoregressive computation into parallel, segment-local prefill verification. Combined with Merkle commitments and signature anchoring, this design achieves strong verifiability and high throughput, ensuring that even a single honest verifier can detect and prove any inference deviation with overwhelming probability [qu2025zkgpt, sun2024zkllm].
5.7 Comparison, Sampling, and On-Chain Adjudication

### v4

@@ 5.2. Registration
  1. (1)
    the scheduler’s VRF verification public key, enabling third parties to validate the pseudorandom outputs used in node assignment and hidden-state sampling (Micali et al., 1999);
  2. (2)
@@
The VRF key ensures that all sampling and scheduling randomness remains publicly auditable and resistant to precomputation or selective bias (Buterin, 2014; Wood, 2014).
@@
This mechanism parallels the delayed-withdrawal model of accountable distributed systems (Castro and Liskov, 1999; Luu et al., 2016), guaranteeing that both schedulers and workers remain economically bound to their past behavior.
5.3. Scheduling Scheme
@@
Verifiable randomness and auditability. Each scheduling decision is deterministically derived from a publicly verifiable random beacon instantiated via a verifiable random function (VRF) (Micali et al., 1999). The Scheduler’s VRF verification key is registered on-chain, and every scheduling round includes the tuple: $(\texttt{job\_id},\texttt{VRF\_out},\texttt{VRF\_proof})$, where VRF_out encodes the pseudorandom seed used for node assignment and hidden-state sampling, and VRF_proof allows any observer or contract to independently validate correctness. This construction guarantees that role assignment and verification checkpoints are unpredictable yet verifiable, preventing any party from biasing job allocation.
@@
(5) $G_{m,\zeta}=\{n\in\mathcal{N}|\texttt{model}(n)=m,~\texttt{slice}(n)=\zeta\},$
@@
(6) $\mathcal{P}(m)=\langle G_{m,1},G_{m,2},\dots,G_{m,L}\rangle.$
@@
(7) $T_{\text{total}}=\sum_{\zeta=1}^{L}\big(T_{\zeta}+T_{\text{compute}}^{(\zeta)}\big),$
@@
5.4. Task Scheduling and Hidden-State Management
@@
(8) $h\leftarrow H(\texttt{req}),\qquad r_{i}\leftarrow\mathrm{VRF}(sk_{\mathsf{sch}},h,i,L_{i}),$
@@
(9) $\mathrm{VRF.Verify}(vk_{\mathsf{sch}},\langle h,i,L_{i}\rangle,r_{i},\pi_{i})=1.$
@@
(10) $n^{\mathrm{inf}}_{i}=G_{m,i}[\sigma_{i}(1)],\qquad\{n^{\mathrm{ver}}_{i,1},\dots,n^{\mathrm{ver}}_{i,k}\}=G_{m,i}[\sigma_{i}(2..k{+}1)].$
@@
(11) $S^{(t)}_{2}=g_{1}\big(\mathrm{Embed}(\mathbf{y}_{<t})\big).$
Each subsequent segment $i=2,\dots,L$ receives the preceding output $S^{(t)}_{i}$, executes its corresponding model layers, and produces: $S^{(t)}_{i+1}=g_{i}(S^{(t)}_{i})$. At the final stage, the tail segment computes the logits and predicts the next token:
(12) $y_{t+1}=\arg\max f_{\mathrm{LM}}\big(S^{(t)}_{L+1}\big)$.
@@
Hidden-State Logging and Commitment. To enable verifiable post-hoc auditing, each Inferencer logs its intermediate hidden states for every token and commits to them using cryptographic proofs. For each token index $t$, an Inferencer serializes its output tensor $S^{(t)}_{i}\in\mathbb{R}^{d_{k}\times d_{v}}$, constructs a Merkle tree $\mathsf{Merkle}(S^{(t)}_{i})$ over all scalar entries, and submits the Merkle root $\gamma^{(t)}_{i}$ accompanied by a bound signature: $\texttt{Sig}_{\mathrm{inf}}\big(h,i,t,\gamma^{(t)}_{i}\big)$. To prevent tampering or substitution during relay, the Scheduler also signs the forwarded root: $\texttt{Sig}_{\mathrm{sch}}\big(h,i,t,\gamma^{(t)}_{i}\big)$, ensuring accountability for both the producer and the forwarder. This dual-signature commitment model provides end-to-end traceability: any discrepancy between a segment’s output and the forwarded record can be attributed with cryptographic certainty. Because verifiers only require Merkle openings for sampled positions, the bandwidth and on-chain verification costs scale sublinearly with tensor size. The integration of deterministic serialization, Merkle-based commitment (Merkle, 1989), and chained signatures provides tamper-evident provenance for all hidden states without introducing measurable latency into the inference path.
Overall, this architecture ensures that every stage of distributed inference remains cryptographically auditable while preserving throughput and latency comparable to optimized large-model serving systems (Rajbhandari et al., 2022; Shoeybi et al., 2019).
5.5. Inference Scheme
Figure 2. Overview of the decentralized inference architecture. Each segment hosts identical Transformer layers, and hidden states are relayed across the pipeline.
The inference process executes online, autoregressive generation over a decentralized, pipelined fabric of homogeneous node groups. Each group hosts an identical copy of a contiguous Transformer block, including weights, normalization statistics, and kernel configurations (Shoeybi et al., 2019; Rajbhandari et al., 2022). For each incoming task, the Scheduler selects eligible groups and, through a verifiable random function (VRF) (Micali et al., 1999), assigns one Inferencer per group to execute the live forward pass. Verifier roles are simultaneously drawn from the same VRF distribution but remain inactive during this phase. Throughout generation, the Scheduler relays and logs hidden states across segment boundaries, appending cryptographic commitments and signatures to construct an auditable trace that serves as the foundation for subsequent verification (Zhang et al., 2020; Chen et al., 2024).
@@ 5.5. Inference Scheme
(13) $S^{(t)}_{1}$ $\leftarrow\texttt{Embed}(x_{1:t}),$
(14) $S^{(t)}_{i+1}$ $\leftarrow F_{i}\big(S^{(t)}_{i};W_{i}\big),\quad i=1,\dots,L,$
(15) $y_{t}$ $\leftarrow\texttt{Decode}\big(S^{(t)}_{L+1}\big)$,
where $S^{(t)}_{i}$ is the hidden state entering segment $i$, $F_{i}$ denotes the segment’s forward operator, $W_{i}$ its parameters, and Decode the final output head combined with the sampling rule. The resulting pipeline achieves token-serial yet segment-parallel execution, synchronized by the Scheduler’s relay of hidden states (Rajbhandari et al., 2022).
Scheduler-Driven Role Assignment. For each inference task, the Scheduler enumerates all node groups capable of collectively serving the model and assigns roles in a publicly verifiable manner. Using a VRF bound to the task identifier (Micali et al., 1999), it deterministically selects one Inferencer per group, while keeping the randomness unpredictable before the VRF proof is revealed. This mechanism ensures that role assignment cannot be biased or precomputed and that every selection remains verifiable through the VRF proof posted on-chain (Bonneau et al., 2015; Qu et al., 2025b).
@@
  • computes a cryptographic commitment to $S^{(t)}_{i+1}$ (e.g., a Merkle root over a deterministic serialization (Merkle, 1989)) and signs the root with its node key; and
@@
The Scheduler verifies each signature against the on-chain registry (optionally through a zero-knowledge membership proof (Chen et al., 2024; Sun et al., 2024) to preserve role anonymity), appends its own relay signature, persists the state to stable storage, and forwards the package to the next segment’s Inferencer. To minimize latency, each receiving node can begin its computation immediately while performing signature verification asynchronously; if any proof later fails, the Scheduler aborts the affected pipeline and invalidates dependent results.
@@
All node public keys are registered under an on-chain Merkle accumulator, enabling compact membership proofs (Micali et al., 1999; Bonneau et al., 2015). Instead of forwarding raw signatures, the Scheduler may alternatively issue a zero-knowledge proof attesting that it holds a valid signature from some registered key on the advertised root—thereby maintaining verifiability without revealing node identity (Sun et al., 2024; Qu et al., 2025b). Each forwarded artifact is re-signed by the Scheduler to guarantee origin authenticity and replay protection. This layered commitment and signature discipline ensures that every hop in the relay chain is both tamper-evident and cryptographically attributable.
Autoregressive Token Emission and Feedback. The terminal segment ($i=L$) outputs logits or the final hidden state for token $t$, from which the next token $y_{t}$ is sampled according to the decoding policy (Brown et al., 2020). The Scheduler collects $y_{t}$ and feeds it back into the pipeline to form $x_{1:t+1}$, advancing the process to step t+1t{+}1. The loop continues until the model emits an end-of-sequence token $\langle\texttt{EOS}\rangle$ or another termination condition is met.
Task-Type Indistinguishability and Robustness. The communication protocol, message format, and API surface used by Inferencers during inference are identical to those used by Verifiers in the subsequent verification phase (Yuan et al., 2022). Consequently, no node can distinguish in real time whether it is performing live inference or participating in a verification probe. This indistinguishability eliminates the opportunity for strategic deviation and simplifies implementation symmetry across both phases. Combined with VRF-driven role assignment, signed commitments, and auditable relays, the inference scheme produces a verifiable, tamper-evident execution trace with negligible online overhead.
5.6. Verification Scheme
Figure 3. Overview of the decentralized verification architecture. Each verifier independently reconstructs its segment’s outputs via a full-sequence prefill.
During autoregressive inference, an LLM executes two phases: a prefill pass—where the prompt tokens are processed in parallel to populate the key–value (KV) cache—and a decode phase—where tokens are generated sequentially with causal dependencies on all previous outputs (Brown et al., 2020; Shoeybi et al., 2019). The decoding phase dominates latency and is inherently serial, leading to limited GPU utilization and memory-bound performance (Rajbhandari et al., 2022). However, verification differs fundamentally: once the full output sequence $y_{1:T}$ is known, the verifier is no longer constrained by sequential dependencies. Instead, it can recompute all token representations simultaneously under the same causal mask, yielding a single parallel pass equivalent to prefill over the concatenated prompt–output sequence (Sun et al., 2024; Zhang et al., 2020).
@@ 5.6. Verification Scheme
(16) $x=[\texttt{prompt}\|y_{1:T}],$
and let $\mathcal{S}=[S^{\text{prompt}}\|S^{1}\|\dots\|S^{T}]$ denote the concatenated hidden states across all positions. Given the causal structure of the Transformer, the verifier can execute a single forward pass to reconstruct all hidden states $\{S^{(t)}_{i}\}_{t=1}^{T}$ for every segment $i\in[1,L]$ without performing incremental token decoding (Vaswani et al., 2023). This transformation replaces $T$ serial decode iterations with a single batched forward computation, eliminating redundant KV-cache updates and reducing cross-node communication. In practice, this yields significantly improved GPU utilization, especially on architectures with high Tensor Core throughput (Rajbhandari et al., 2022).
@@
(17) $T_{\text{verify}}\approx\frac{1}{L}T_{\text{prefill}}(\text{full model}),$
yielding near-linear scaling in the number of participating verifier groups (Rajbhandari et al., 2022; Zhang et al., 2020).
Figure 4. Verification workflow. Each verifier independently executes a full-sequence prefill over its model segment and commits its output root to the blockchain for later consistency checks.
Prefill Computation. After the inference phase concludes, the Scheduler initiates segment-local verification by distributing the required input states to each verifier (Qu et al., 2025b). For segment $i>1$, the Scheduler sends the sequence $\{S_{i}^{(t)}\}_{t=1}^{T}$—the hidden states produced by the upstream inferencer for all tokens. For the first segment ($i=1$), the Scheduler instead provides the full tokenized input $\texttt{prompt}\|y_{1:T}]$, corresponding to the raw embedding input of the model. Upon receipt, each verifier executes a full-sequence prefill limited to its hosted layers:
(18) $S_{i+1}^{(t)}=F_{i}(S_{i}^{(t)};W_{i}),\quad\forall t\in[1,T],$
@@
Importantly, the communication interface, data serialization format, and RPC protocol used during verification are identical to those used in live inference (Yuan et al., 2022). This design choice enforces task-type indistinguishability: from the perspective of any node, it is impossible to discern whether the current request corresponds to live inference or a verification probe (Bonneau et al., 2015). Such indistinguishability eliminates mode-specific behavior and prevents nodes from selectively deviating or shortcutting computation when acting as verifiers. Upon completion, each verifier must return its computed outputs $\{S^{(t)}_{i+1}\}_{t=1}^{T}$ through the same interface as during inference, producing an auditable trace indistinguishable in structure and timing.
Commitment Submission. Once the local recomputation completes, the Scheduler transitions verifiers into the commit phase. Each verifier constructs a tamper-evident cryptographic commitment to its final outputs—specifically, the hidden state corresponding to the last token ($t=T$) at its segment’s output boundary. This commitment is instantiated as a Merkle root (Merkle, 1989):
(19) $\texttt{root}^{(T)}_{i}=\texttt{MerkleRoot}\big(S^{(T)}_{i+1}\big)$,
@@
(20) $\sigma^{(T)}_{i}=\texttt{Sig}(\texttt{sk}_{i},\texttt{root}^{(T)}_{i})$,
@@
and publishes the commitment on-chain through the designated verification contract (Chen et al., 2024; Sun et al., 2024).
@@
This on-chain record provides a permanent anchor for later verification steps. In subsequent sampling or dispute phases, verifiers may be required to disclose specific slices or elements of the hidden state $\{S^{(T)}_{i+1}\}$, accompanied by Merkle inclusion proofs that can be efficiently checked against the posted root (Bonneau et al., 2015; Micali et al., 1999). This approach ensures both compactness and auditability: the entire verification process involves only lightweight commitments and selective disclosures rather than full tensor uploads, keeping blockchain overhead minimal while preserving cryptographic integrity.
@@
In summary, the verification scheme leverages the known output sequence and the pre-recorded intermediate states to convert sequential autoregressive computation into parallel, segment-local prefill verification. Combined with Merkle commitments and signature anchoring, this design achieves strong verifiability and high throughput, ensuring that even a single honest verifier can detect and prove any inference deviation with overwhelming probability (Qu et al., 2025b; Sun et al., 2024).
5.7. Comparison, Sampling, and On-Chain Adjudication